OpenStudy (anonymous):

Write a formula for this sequence: 1/5, 3/5, 1, 1 2/5..

OpenStudy (kainui):

Hmm, is there some way we can try to find out the difference between what happens at each step to see if there's some pattern there?

OpenStudy (anonymous):

thats all i have. /:

OpenStudy (kainui):

How much did they add between the first number and the second number?

OpenStudy (anonymous):

2

OpenStudy (kainui):

Good! Although don't forget those 5's at the bottom, it doesn't really change much except they added 2/5.

OpenStudy (kainui):

So how much did they add to get from the second number to the third number?

OpenStudy (anonymous):

2/5?

OpenStudy (kainui):

Yeah! After all, 3 fifths added to 2 fifths is 5 fifths, which makes sense that would be 1. Alright, now let's see if we can make a formula. Try some stuff out and don't be afraid to fail!

OpenStudy (anonymous):

but 2/5's added to 1 doesnt match with 2/5's

OpenStudy (kainui):

The last number is the mixed number \[\Huge 1 \frac{2}{5}\] correct?

OpenStudy (anonymous):

i did not see the 1 infront, but yes it is

OpenStudy (kainui):

Ahh, well now that you see it, it makes sense right?

OpenStudy (anonymous):

no.

OpenStudy (kainui):

What term is not 2/5 greater than the previous term?

OpenStudy (anonymous):

wait yes it makes sense

OpenStudy (kainui):

Ok cool. So now every term is really just the first term with a bunch of extra 2/5 added to it right? We could rewrite the sequence as: \[\Large \frac{1}{5},\ \ \frac{1}{5}+\frac{2}{5}, \ \ \frac{1}{5}+\frac{2}{5}+\frac{2}{5}, \ \frac{1}{5}+\frac{2}{5}+\frac{2}{5}+\frac{2}{5}, \ \ etc\] See what I mean? But is there a better way to do this?

OpenStudy (anonymous):

im not sure

OpenStudy (kainui):

If I give you 3+3+3+3 would you say that's the same as 4*3 since there are 4 threes?

OpenStudy (anonymous):

yes

OpenStudy (kainui):

Let's look at a slightly different sequence then: 1, 4, 7, 10 Notice how we can rewrite this as 1, 1+3, 1+3+3, 1+3+3+3 since each term is just 3 greater than the last But now notice we can combine terms to get: 1+0*3, 1+1*3, 1+2*3, 1+3*3 So if we call each term by a number then we can say that the first term is 0, the next term is 1, the next is 2, etc... then we can write down that 1+3n will be the formula for this sequence. So if I want the term where n=3 then I plug it in and get 1+3*3=10 and if we want the term before n=3 we plug in n=2 1+2*3=7 and sure enough we are getting the same sequence!

OpenStudy (anonymous):

im quite confused

OpenStudy (kainui):

By what?